# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ III-B3 ZKP based Verifiable Inference
In 2020, Froelicher et al. [35] proposed a decentralized ZKML system named Drynx that combines interactive protocols, homomorphic encryption, zero-knowledge proofs, and differential privacy technologies, enabling privacy-preserving statistical queries (inferences) and training and evaluation of logistic regression models on distributed datasets. The Drynx system consists of four main components: the querier (Q), data providers (DPs), computation nodes (CNs), and verification nodes (VNs). Within the system’s interaction protocol, Q initiates a query request, after which the DPs encrypt their raw data using homomorphic encryption techniques and transmit the encrypted responses to the CNs. The CNs aggregate and process the received encrypted data collectively, generating both the query results and corresponding zero-knowledge proofs to ensure the correctness of the computation process. The VNs are responsible for verifying the correctness of the computations. Through this interaction protocol, Drynx effectively guarantees the transparency and verifiability of query execution. Even under a strong adversarial model with malicious entities, the system ensures the correctness of the computation results. Drynx employs Camenisch-Stadler ZKP [58] to verify the computational integrity of CNs, and Camenisch-Chaabouni ZKP [59] to validate the range of input data, ensuring that the data provided by DPs falls within legitimate bounds. CNs generate zero-knowledge proofs upon completing calculations, while VNs collectively verify these proofs to ensure the accuracy of computational results, with verification results stored on the blockchain to achieve computational process auditability. Additionally, the Drynx system implements differential privacy protection based on the Collective Differential Privacy (CDP) protocol introduced in Unlynx [60], effectively protecting individual privacy by adding noise to query results.
@@
vCNN [37] is a verifiable inference framework specifically designed to accelerate the proof process for convolutional neural networks (CNNs). In CNNs, the output of a convolutional operation is represented as a sum of products: $y_{i}=\sum\nolimits_{j=1}^{l-0}{a_{j}\cdot x_{l-1-j}}$, where $y_{i}$ is the $i$-th convolutional output, $\left(x_{i},\cdots,x_{i+l-1}\right)$ is the $i$-th input data vector of length $l$, and $\left(a_{0},\cdots,a_{n-1}\right)$ is the convolutional kernel vector of length $n$. The number of multiplications involved in a convolutional operation is $O(ln)$, which is often computationally expensive. Consequently, traditional zk-SNARKs incur significant proving time when verifying these multiplications. To address this challenge, vCNN introduces a novel approach to reduce the cost of proving CNN convolutional operations by transforming the sum of products into a product of sums: $\left(\sum\nolimits_{i=1}^{n+l-2}y_{i}\right)=\left(\sum\nolimits_{i=0}^{n-1}x_{i}\right)\left(\sum\nolimits_{i=0}^{l-1}a_{i}\right)$. However, this transformation introduces a potential issue: multiple different values of $y_{i}^{\prime}\neq y_{i}$ can satisfy the above equality if $\sum\nolimits_{i=1}^{n+l-2}y_{i}^{\prime}=\sum\nolimits_{i=1}^{n+l-2}y_{i}$. To prevent such ambiguities, an indeterminate variable $Z$ is introduced, ensuring that the following equality holds for all $Z$: $\left(\sum\nolimits_{i=1}^{n+l-2}y_{i}Z^{i}\right)=\left(\sum\nolimits_{i=0}^{n-1}x_{i}Z^{i}\right)\left(\sum\nolimits_{i=0}^{l-1}a_{i}Z^{i}\right)$. This equality corresponds to a polynomial multiplication: $y(Z)=x(Z)\cdot a(Z)$, where $y(Z)$, $x(Z)$, and $a(Z)$ are polynomials. The verifiability of this transformation can be encoded using quadratic polynomial programs (QPPs) [61], reducing the proving cost from $O(ln)$ to $O(n+l)$ when $n$ and $l$ take practical values. This reduction in complexity enables vCNN to significantly improve the proving time for CNN inference compared to traditional zk-SNARKs based on quadratic arithmetic programs (QAPs) [62]. Experimental results demonstrate the effectiveness of vCNN. On the simple MNIST dataset, vCNN improves proof performance by a factor of 20. For the more complex VGG16 model [63], it achieves an improvement of 18,000 times.
Due to the lack of efficient ZKP protocols capable of verifying the inference results produced by the complex computations of neural networks, Weng et al. [13] proposed Mystique. This system is built upon the subfield vector oblivious linear evaluation (sVOLE) [64] interactive proof protocol. The work introduces three main contributions. (1) Mystique provides a set of efficient ZKP building blocks tailored for implementing inference in large-scale neural networks. These building blocks enable developers to construct ZKML systems with ease, abstracting away the complexities of the underlying cryptographic logic. (2) Three optimized ZKP protocols are proposed to address various challenges: (i) A novel protocol for efficient conversion between arithmetic and Boolean values, which enhances performance by adapting the circuit to specific computational requirements; (ii) A protocol for efficient conversion between public commitments and private authenticated values, facilitating the seamless integration of publicly committed values into ZKP systems; (iii) A protocol for efficient conversion between fixed-point and floating-point numbers, resolving the inconsistency between floating-point representations used in machine learning algorithms and fixed-point representations used in cryptographic computations. (3) Mystique introduces an optimized zero-knowledge proof protocol for matrix multiplication, achieving sublinear private multiplications relative to the size of the matrix. This significantly improves computational efficiency for matrix-related operations in ZKP systems.
@@
Singh et al. proposed a zero-knowledge verifiable scheme for a distributed AI pipeline, which includes a privacy-preserving verification protocol for the inference using decision trees [40]. The distributed AI pipeline assigns different stages of the process—such as data collection, model training, and model inference—to independent participants, including data owners, model owners, and model users. However, previous approaches faced significant inefficiencies in handling memory within the circuit. To overcome this limitation, the authors introduced an improved protocol for the privacy-preserving verifiable inference using decision trees. By eliminating the need for expensive one-time hash operations on the tree structure, the size of the verification circuit was reduced by up to tenfold. Additionally, the protocol leverages read-only memory access to further optimize performance, significantly reducing the number of multiplication gates required per prediction. Experimental results demonstrate that, compared to zkDT—a similar scheme designed for the verification of decision tree inference—the proposed protocol achieves substantial improvements in efficiency, with its circuit size being only $25\%$ to $40\%$ of zkDT.
@@
In December 2022, Wang et al. proposed ezDPS [11], an efficient machine learning inference pipeline (MLIP) designed to process data across multiple stages at the ML service provider’s end. The pipeline enables the service provider, who supplies the ML model, to compute the final inference result for the client, who provides the input data, while safeguarding the private model parameters at each stage. After committing to its private model, the ezDPS MLIP operates through three key processing stages: (i) utilizing discrete wavelet transform (DWT) for initial data transformation; (ii) applying principal component analysis (PCA) to reduce dimensionality and extract essential features; (iii) performing classification using support vector machines (SVM). The framework employs Hyrax [65] as the underlying polynomial commitment scheme and Spartan [66] as the backend zero-knowledge proof (ZKP) protocol to verify computations across these stages. Furthermore, ezDPS introduces the concept of zero-knowledge proof of accuracy (zkPoA), which allows the ML service provider to prove the accuracy of the committed model on public datasets without revealing the model parameters. However, a notable limitation of this approach is its inability to preserve client data privacy. Specifically, clients must send their input data in plaintext to the ezDPS platform for computation, potentially exposing sensitive information.
@@
Kang et al. [39] leveraged the Halo2 ZKP scheme [67] to construct a zk-SNARK arithmetic circuit capable of supporting the MobileNet v2 model [56]. In contrast to prior works that primarily targeted small-scale datasets (e.g., MNIST or CIFAR-10) with simpler models, this study demonstrates that zk-SNARKs can effectively handle real-world, large-scale neural network models. Specifically, it supports MobileNet v2, a model trained and evaluated on the large-scale ImageNet dataset—a widely used computer vision benchmark comprising over 1 million high-resolution images across 1,000 categories. This work primarily addresses the computational overhead of verifying division operations in the circuit, which is a significant bottleneck in zk-SNARK-based deep learning inference. To tackle this, the authors proposed two key optimizations based on the Plonkish arithmetization framework [68] in the Halo2 scheme. (1) For linear operations, including convolutional layers, residual connection layers, and fully connected layers, the authors designed two custom gates to efficiently encode division operations. The first gate performs an addition of multiple inputs, while the second gate computes the dot product of inputs and weights, incorporating the division by a fixed scaling factor. These gates avoid the need for complex floating-point arithmetic by leveraging fixed-point approximations, significantly reducing the number of constraints required for division. (2) For nonlinear operations such as ReLU and softmax, the authors utilized lookup arguments to efficiently represent division. Specifically, the division operation is precomputed for a range of possible input values, and the results are stored in a lookup table. During proof generation, the circuit enforces that the computed division results match the precomputed values in the table, thereby reducing the computational cost associated with division. Additionally, the authors ensured the privacy of the model’s inputs and weights by incorporating hash commitments. This allows the model provider to prove the correctness of inference without revealing sensitive model parameters or input data. The proposed zk-SNARK scheme achieves high accuracy on ImageNet-scale models while maintaining relatively low verification time. For example, the proof verification time for MobileNet v2 with 79.2% accuracy on ImageNet is approximately 10.27 seconds on commodity hardware. Furthermore, the proof size is significantly smaller compared to prior methods. Specifically, the proof size is reduced to 5,952 bytes, which is orders of magnitude smaller than methods based on secure multi-party computation (MPC), which typically require tens to hundreds of gigabytes for proof representation.
@@
Fan et al. proposed a zero-knowledge verification scheme to ensure the integrity of CNN inference, with a focus on balancing the confidentiality of certain CNN model parameters and verifying the correctness of the inference process in machine learning as a service (MLaaS) scenarios [42]. In this scheme, the authors developed a computational logic extraction algorithm capable of accurately translating the computational logic of convolutional layers, fully connected layers, pooling layers, and activation layers into corresponding simple arithmetic expressions. These expressions are subsequently used to construct proof circuits for each layer of the model within the zk-SNARK framework. Based on this design, the prover performs the computations involved in the CNN model’s inference process and generates zero-knowledge proofs using the constructed proof circuits. The verifier then validates the integrity of the inference process by checking the proofs for each layer of the model. Although this scheme increases the number of multiplications and additions, resulting in a more complex R1CS for circuit construction, experimental results show that both the computational and storage overheads remain within acceptable limits. This demonstrates the practicality of the proposed approach for real-world applications.
Ganescu et al. [43] proposed a novel ZKML scheme, named snarkGPT, to address the challenge of securely executing Generative AI models via remote cloud APIs for all users, without exposing sensitive model information. Since transformers serve as the foundational architecture for many Generative AI models, snarkGPT is specifically designed to prove the execution of transformer-based models. Using a fixed version of the EZKL framework (commit 8f122bf1), Ganescu et al. implement the snarkGPT protocol based on Halo2, with a particular focus on the nanoGPT model [69]. The performance of their implementation is evaluated in terms of time and memory costs associated with proof generation, particularly for the prover, who is responsible for generating the zk-SNARK proof. Experimental results show that proof generation time increases nonlinearly with both embedding size and layer count. Additionally, the study examines the impact of circuit matrix size on proof generation time and memory consumption, revealing that larger matrix sizes lead to significant increases in both. Specifically, when the logarithmic size of the circuit matrix grows from 14 to 18, runtime improves slightly, but as the size reaches 24 and 25, there is a sharp increase in both runtime and memory costs. Notably, memory usage escalates dramatically, reaching 148 GB for a logarithmic matrix size of 25. The research also explores the relationship between model architecture, the number of parameters, and the number of constraints generated in zk-SNARK proofs. Experiments highlight a stark contrast in the constraint-to-parameter ratio between nanoGPT and other models, such as the Modulus Labs MLP model. For nanoGPT, this ratio ranges from approximately 58x to 85x, whereas the MLP model exhibits a significantly lower ratio. By comparing constraint generation across different models, the study uncovers substantial differences in zk-SNARK proof generation between Transformer architectures (like nanoGPT) and other types of neural networks, such as Convolutional Neural Networks. These findings suggest that zk-SNARK proof generation time and memory costs are influenced not only by the number of model parameters but also by factors such as architectural design and the choice of proof system.
Handling large-scale machine learning models requires substantial memory and computational resources, which has historically constrained previous ZKML schemes to relatively small-scale models. To overcome this limitation, Chen et al. proposed a novel system capable of generating zero-knowledge proofs for the inference of larger-scale machine learning models, including contemporary vision models and a lightweight version of GPT-2 [45]. This system addresses the challenge of proving large-scale models by compiling models built in TensorFlow into circuits within the Halo2 proof system. The core contribution lies in the introduction of an optimized compiler, which features efficient constraint gadgets for fundamental operations and an advanced optimizer to determine the optimal layout for these gadgets. This innovation dramatically improves the compilation perfoHrmance, achieving up to 24$\times$ speedups for implementing the same operations in zk-SNARK circuits. In comparison to prior works, the combination of the optimized compiler and the Halo2 framework enables the system to support a broader range of models while delivering significant improvements in proof generation speed, verification time, and proof size. This advancement represents a major step forward in making zero-knowledge proofs practical for large-scale machine learning applications.
zkLLM [46] is a solution for verifiable inference of large language models (LLMs) based on the transformer architecture, aiming to verify the authenticity of LLM inference outputs without disclosing the model parameters. Its implementation can be summarized into the following two key components. (1) tlookup: The paper proposed the tlookup protocol to address the high computational complexity and memory overhead associated with non-arithmetic tensor operations (such as activation functions) in ZKP. The tlookup protocol pre-constructs a lookup table that stores the mapping between activation function inputs and outputs. This approach eliminates the need to directly compute complex nonlinear functions during the proof process. Instead, it enables rapid verification of input-output pairs through table lookup, thereby significantly reducing computational complexity. Additionally, tlookup introduces a random linear combination technique to compress input and output values, reducing the amount of data that needs to be transmitted and processed during proof generation. To further enhance efficiency, the tlookup protocol fully leverages modern hardware capabilities (e.g., GPUs) by designing a verification process that supports parallel computation. In this process, the kernel functions for table lookup and random linear combination are parallelized, greatly reducing the time required for inference proof generation. (2) zkAttn: Building upon tlookup, the paper further introduced the zkAttn protocol, which provides an efficient zero-knowledge proof process for the attention mechanism in LLMs. The core computations of the attention mechanism include matrix multiplication and the Softmax function. These operations typically incur high computational costs and storage overhead in zero-knowledge proofs. zkAttn introduces specific optimization methods, such as exploiting the shift invariance of the Softmax function and applying simplified and piecewise linear approximations to the input-output pairs of the Softmax function. These strategies significantly reduce the complexity of proof generation. Through these designs, even the most complex attention mechanism layers in LLMs can be efficiently subjected to zero-knowledge verification. Through these two components, zkLLM provides a fully parallelized scheme based on CUDA technology for verifying the correctness of large language model inference results without exposing the model’s private parameters. Experimental results demonstrate that zkLLM exhibits efficient performance on large language models with up to 13 billion parameters (such as OPT and LLaMa-2). Specifically, zkLLM’s proof generation time is under 15 minutes, and verification time requires only a few seconds, significantly enhancing verification efficiency. Furthermore, the scheme keeps GPU memory consumption within 23.1 GB, making it compatible with common GPU hardware environments. In terms of numerical accuracy, zkLLM performs well, achieving performance close to the original model and ensuring the correctness of inference results.
ZKML [45] is a compiler specifically designed to optimize zero-knowledge proofs for machine learning models. Its core functionality involves converting TensorFlow models into Halo2 zero-knowledge proof circuits, enabling efficient verification of complex machine learning inference processes. The system architecture comprises two critical modules: low-level gadgets and a Circuit Layout Optimizer. First, the low-level gadgets serve as fundamental components of ZKML, providing efficient constraint representations for common machine learning operations such as dot products, Softmax, and nonlinear activation functions (e.g., ReLU, GELU). These constraints enable the implementation of these operations within zero-knowledge proof circuits with minimal computational overhead and storage requirements, thereby supporting diverse model architectures including GPT-2, Twitter models, ResNet-18, and VGG-16. Second, the Circuit Layout Optimizer automatically selects optimal circuit configurations through circuit simulation and cost estimation. This intelligent optimization module adapts circuit structures based on hardware resources and model characteristics, effectively reducing computational redundancy and significantly improving proof efficiency. Through these dual mechanisms, ZKML demonstrates remarkable performance improvements and flexibility in supporting zero-knowledge proofs for large-scale machine learning model inference. Experimental evaluations on multiple models (including GPT-2, Diffusion Models, Twitter models, ResNet-18, and VGG-16) reveal compelling results: For GPT-2, proof generation requires approximately 3,652 seconds with verification completed in 18.7 seconds and a compact proof size of 28 KB. ResNet-18 achieves proof generation in 52.9 seconds, verification in 12 milliseconds, and a proof size of 15.3 KB. Comparative analysis with existing zero-knowledge proof systems (zkCNN [10] and vCNN [37]) demonstrates ZKML’s superior performance, achieving up to 24-fold improvement in proof generation time, 5x reduction in verification time, and 22x compression in proof size. Notably, on the CIFAR-10 dataset, ZKML reduces proof generation time by 31 hours compared to conventional approaches. Furthermore, the system maintains exceptional model accuracy with negligible impact (within 0.01%), significantly outperforming traditional quantization methods that typically incur 0.1% accuracy degradation, thereby achieving an optimal balance between efficiency and precision.
Lu et al. [47] proposed an efficient and scalable ZKP framework aimed at optimizing the generation of verifiable inference proofs for neural networks. By integrating the Vector Oblivious Linear Evaluation (VOLE) [70] technology, range proofs, and lookup proofs, this framework significantly improves the proof efficiency of non-linear layers and can be extended to more types of neural networks. The framework not only supports CNNs, such as ResNet-101 and VGG-11, but also Transformer networks like GPT-2, supporting Transformer networks with up to 117 million parameters. The specific details of the framework proposed in [47] are as follows. (1) To tackle the issue of high computational overhead in the non-linear layers of neural networks, the framework transforms complex non-linear relationships into two core proof forms: Range Proofs and Exponent Proofs. This approach significantly reduces the large number of constraints that traditional methods need to handle. Range Proofs are used to verify whether input values fall within a specific range, while Exponent Proofs deal with complex relationships involving exponential operations, effectively reducing computational complexity. Simultaneously, based on VOLE technology, the framework designs two improved methods: a Range Proof module and a Lookup Proof. These methods not only efficiently perform range verification but also rapidly complete data lookup and mapping operations, greatly enhancing the efficiency and scalability of proof generation. (2) The paper adopts a modular design approach, endowing the framework with high flexibility and scalability. Through this modular method, the framework can incrementally combine basic operational units (such as addition, multiplication, and range verification) to extend support to the inference verification of entire neural networks. Specifically, the framework optimizes CNNs and Transformer networks, enabling efficient proof and verification of the inference processes and results of these common models. Additionally, this modular design provides the possibility for the framework to extend to other types of neural networks (such as Recurrent Neural Networks or Graph Neural Networks), allowing it to maintain efficiency and generality across different application scenarios. Through these two key steps, the paper provides an efficient and scalable neural network verification scheme while ensuring data privacy protection. Experimental results demonstrate significant performance improvements in non-linear layer proofs and overall neural network verification. In the proofs of non-linear layers (such as ReLU, Softmax, GELU, and Normalization), compared to Mystique (a scheme based on bit decomposition), the proof generation time achieved up to a 477.2x speedup, and the computational and communication costs were also significantly better than existing schemes. In the neural network inference proof process, significant optimizations were achieved for Convolutional Neural Networks (such as VGG-11, ResNet-50, and ResNet-101). For example, the proof generation time for ResNet-101 was only 6.65 seconds, an acceleration of 41.4 times compared to traditional methods; for Transformer Neural Networks (such as GPT-2), the proof generation time was 287.1 seconds, a 35.7x improvement over ZKML [46]. Furthermore, the framework is compatible with high-precision quantized neural networks, achieving inference accuracy close to the original model. For instance, the accuracy of ResNet-101 on the CIFAR-10 dataset decreased by only 0.04% , fully demonstrating its efficiency and reliability.

### v2

@@ III-B3 ZKP-based Verifiable Inference
In 2020, [37] proposed a decentralized ZKML system named Drynx that combines interactive protocols, homomorphic encryption, zero-knowledge proofs, and differential privacy technologies, enabling privacy-preserving statistical queries (inferences) and training and evaluation of logistic regression models on distributed datasets. The Drynx system consists of four main components: the querier (Q), data providers (DPs), computation nodes (CNs), and verification nodes (VNs). Within the system’s interaction protocol, Q initiates a query request, after which the DPs encrypt their raw data using homomorphic encryption techniques and transmit the encrypted responses to the CNs. The CNs aggregate and process the received encrypted data collectively, generating both the query results and corresponding zero-knowledge proofs to ensure the correctness of the computation process. The VNs are responsible for verifying the correctness of the computations. Through this interaction protocol, Drynx effectively guarantees the transparency and verifiability of query execution. Even under a strong adversarial model with malicious entities, the system ensures the correctness of the computation results. Drynx employs Camenisch-Stadler ZKP [59] to verify the computational integrity of CNs, and Camenisch-Chaabouni ZKP [60] to validate the range of input data, ensuring that the data provided by DPs falls within legitimate bounds. CNs generate zero-knowledge proofs upon completing calculations, while VNs collectively verify these proofs to ensure the accuracy of computational results, with verification results stored on the blockchain to achieve computational process auditability. Additionally, the Drynx system implements differential privacy protection based on the Collective Differential Privacy (CDP) protocol introduced in Unlynx [61], effectively protecting individual privacy by adding noise to query results.
@@
vCNN is a verifiable inference framework specifically designed to accelerate the proof process for convolutional neural networks (CNNs) [38]. In CNNs, the output of a convolutional operation is represented as a sum of products: $y_{i}=\sum\nolimits_{j=1}^{l-0}{a_{j}\cdot x_{l-1-j}}$, where $y_{i}$ is the $i$-th convolutional output, $\left(x_{i},\cdots,x_{i+l-1}\right)$ is the $i$-th input data vector of length $l$, and $\left(a_{0},\cdots,a_{n-1}\right)$ is the convolutional kernel vector of length $n$. The number of multiplications involved in a convolutional operation is $O(ln)$, which is often computationally expensive. Consequently, traditional zk-SNARKs incur significant proving time when verifying these multiplications. To address this challenge, vCNN introduces a novel approach to reduce the cost of proving CNN convolutional operations by transforming the sum of products into a product of sums: $\left(\sum\nolimits_{i=1}^{n+l-2}y_{i}\right)=\left(\sum\nolimits_{i=0}^{n-1}x_{i}\right)\left(\sum\nolimits_{i=0}^{l-1}a_{i}\right)$. However, this transformation introduces a potential issue: multiple different values of $y_{i}^{\prime}\neq y_{i}$ can satisfy the above equality if $\sum\nolimits_{i=1}^{n+l-2}y_{i}^{\prime}=\sum\nolimits_{i=1}^{n+l-2}y_{i}$. To prevent such ambiguities, an indeterminate variable $Z$ is introduced, ensuring that the following equality holds for all $Z$: $\left(\sum\nolimits_{i=1}^{n+l-2}y_{i}Z^{i}\right)=\left(\sum\nolimits_{i=0}^{n-1}x_{i}Z^{i}\right)\left(\sum\nolimits_{i=0}^{l-1}a_{i}Z^{i}\right)$. This equality corresponds to a polynomial multiplication: $y(Z)=x(Z)\cdot a(Z)$, where $y(Z)$, $x(Z)$, and $a(Z)$ are polynomials. The verifiability of this transformation can be encoded using quadratic polynomial programs (QPPs) [62], reducing the proving cost from $O(ln)$ to $O(n+l)$ when $n$ and $l$ take practical values. This reduction in complexity enables vCNN to significantly improve the proving time for CNN inference compared to traditional zk-SNARKs based on quadratic arithmetic programs (QAPs) [63]. Experimental results demonstrate the effectiveness of vCNN. On the simple MNIST dataset, vCNN improves proof performance by a factor of 20. For the more complex VGG16 model [64], it achieves an improvement of 18,000 times.
As researchers moved toward larger neural networks, the lack of efficient proof primitives for complex inference workloads became a major bottleneck. Mystique [39] addresses this issue by building an sVOLE-based [65] framework that provides reusable proof components for neural-network inference. Its main contributions are threefold. First, Mystique offers a collection of practical ZKML building blocks that abstract away much of the low-level cryptographic complexity. Second, it introduces efficient conversion protocols between arithmetic and Boolean values, between public commitments and private authenticated values, and between fixed-point and floating-point representations, all of which are central to bridging ML computations with proof-system arithmetic. Third, it presents an optimized matrix-multiplication proof whose number of private multiplications is sublinear in the matrix size. Taken together, these ideas make Mystique an important systems contribution, especially for later work on more expressive and scalable inference verification.
MSCProof [50] pushes this modular direction further with a framework for sequential verifiable computation based on Verifiable Evaluation schemes on Fingerprinted Data (VEs), an abstraction that captures diverse sumcheck-style proofs including GKR. It also designs a VE specialized for batched multi-channel convolution and uses it to build modular proofs for CNN and image-processing pipelines. This work is particularly relevant to ZKML because it narrows the gap between highly tailored ML proof systems and more composable proof-engineering workflows.
@@
[40] proposed a zero-knowledge verifiable scheme for a distributed AI pipeline, which includes a privacy-preserving verification protocol for inference using decision trees. The distributed AI pipeline assigns different stages of the process—such as data collection, model training, and model inference—to independent participants, including data owners, model owners, and model users. However, previous approaches faced significant inefficiencies in handling memory within the circuit. To overcome this limitation, the authors introduced an improved protocol for the privacy-preserving verifiable inference using decision trees. By eliminating the need for expensive one-time hash operations on the tree structure, the size of the verification circuit was reduced by up to tenfold. Additionally, the protocol leverages read-only memory access to further optimize performance, significantly reducing the number of multiplication gates required per prediction. Experimental results demonstrate that, compared to zkDT—a similar scheme designed for the verification of decision tree inference—the proposed protocol achieves substantial efficiency improvements, with its circuit size being only $25\%$ to $40\%$ of zkDT.
@@
In December 2022, [48] proposed ezDPS, an efficient machine learning inference pipeline (MLIP) designed to process data across multiple stages at the ML service provider’s end. The pipeline enables the service provider, who supplies the ML model, to compute the final inference result for the client, who provides the input data, while safeguarding the private model parameters at each stage. After committing to its private model, the ezDPS MLIP operates through three key processing stages: (i) utilizing discrete wavelet transform (DWT) for initial data transformation; (ii) applying principal component analysis (PCA) to reduce dimensionality and extract essential features; (iii) performing classification using support vector machines (SVM). The framework employs Hyrax [66] as the underlying polynomial commitment scheme and Spartan [67] as the backend zero-knowledge proof (ZKP) protocol to verify computations across these stages. Furthermore, ezDPS introduces the concept of zero-knowledge proof of accuracy (zkPoA), which allows the ML service provider to prove the accuracy of the committed model on public datasets without revealing the model parameters. However, a notable limitation of this approach is its inability to preserve client data privacy. Specifically, clients must send their input data in plaintext to the ezDPS platform for computation, potentially exposing sensitive information.
@@
[68] leveraged the Halo2 ZKP scheme to construct a zk-SNARK arithmetic circuit capable of supporting the MobileNet v2 model [69] in the paper [47]. In contrast to prior works that primarily targeted small-scale datasets (e.g., MNIST or CIFAR-10) with simpler models, this study demonstrates that zk-SNARKs can effectively handle real-world, large-scale neural network models. Specifically, it supports MobileNet v2, a model trained and evaluated on the large-scale ImageNet dataset—a widely used computer vision benchmark comprising over 1 million high-resolution images across 1,000 categories. This work primarily addresses the computational overhead of verifying division operations in the circuit, which is a significant bottleneck in zk-SNARK-based deep learning inference. To tackle this, the authors proposed two key optimizations based on the Plonkish arithmetization framework [70] in the Halo2 scheme. (1) For linear operations, including convolutional layers, residual connection layers, and fully connected layers, the authors designed two custom gates to efficiently encode division operations. The first gate performs an addition of multiple inputs, while the second gate computes the dot product of inputs and weights, incorporating the division by a fixed scaling factor. These gates avoid the need for complex floating-point arithmetic by leveraging fixed-point approximations, significantly reducing the number of constraints required for division. (2) For nonlinear operations such as ReLU and softmax, the authors utilized lookup arguments to efficiently represent division. Specifically, the division operation is precomputed for a range of possible input values, and the results are stored in a lookup table. During proof generation, the circuit enforces that the computed division results match the precomputed values in the table, thereby reducing the computational cost associated with division. Additionally, the authors ensured the privacy of the model’s inputs and weights by incorporating hash commitments. This allows the model provider to prove the correctness of inference without revealing sensitive model parameters or input data. The proposed zk-SNARK scheme achieves high accuracy on ImageNet-scale models while maintaining relatively low verification time. For example, the proof verification time for MobileNet v2 with 79.2% accuracy on ImageNet is approximately 10.27 seconds on commodity hardware. Furthermore, the proof size is significantly smaller compared to prior methods. Specifically, the proof size is reduced to 5,952 bytes, which is orders of magnitude smaller than methods based on secure multi-party computation (MPC), which typically require tens to hundreds of gigabytes for proof representation.
@@
[49] proposed a zero-knowledge verification scheme to ensure the integrity of CNN inference, with a focus on balancing the confidentiality of certain CNN model parameters and verifying the correctness of the inference process in machine learning as a service (MLaaS) scenarios. In this scheme, the authors developed a computational logic extraction algorithm capable of accurately translating the computational logic of convolutional layers, fully connected layers, pooling layers, and activation layers into corresponding simple arithmetic expressions. These expressions are subsequently used to construct proof circuits for each layer of the model within the zk-SNARK framework. Based on this design, the prover performs the computations involved in the CNN model’s inference process and generates zero-knowledge proofs using the constructed proof circuits. The verifier then validates the integrity of the inference process by checking the proofs for each layer of the model. Although this scheme increases the number of multiplications and additions, resulting in a more complex R1CS for circuit construction, experimental results show that both the computational and storage overheads remain within acceptable limits. This demonstrates the practicality of the proposed approach for real-world applications.
[51] studied Transformer-scale proving through snarkGPT, an implementation built on Halo2 and nanoGPT [71]. Rather than treating LLM proving as a straightforward extension of CNN-style pipelines, the paper highlights how proving cost grows rapidly with layer width, depth, and circuit matrix size, with memory quickly becoming a dominant constraint. Its main value in the survey is therefore diagnostic: it makes clear that Transformer verification requires different engineering assumptions from earlier small-model ZKML systems.
[41] moves closer to large-model practical deployment by introducing a Halo2-based compiler for TensorFlow models together with efficient low-level gadgets and a circuit-layout optimizer. The system supports a broader range of architectures, including vision models and smaller Transformer variants, while improving proving time, verification time, and proof size relative to earlier frameworks. More importantly, it shows that compiler quality and layout optimization are now as important as proof-system choice for practical ZKML inference.
zkLLM [42] extends verifiable inference to large language models through two main components: tlookup for lookup-based handling of nonlinear tensor operations, and zkAttn for more efficient verification of attention and Softmax-related computation. By exploiting lookup compression and GPU parallelism, it demonstrates that LLM inference proofs can be generated for models up to 13B parameters within practical memory and time budgets. This makes zkLLM a milestone in pushing ZKML beyond CNN-scale inference.
Lu et al. [43] propose a complementary inference framework built around VOLE, range proofs, and lookup proofs. Their design focuses especially on reducing the cost of nonlinear layers and keeping the system modular enough to support both CNNs and Transformer models. Compared with earlier bit-decomposition-heavy approaches, this line of work shows that alternative proving backends can yield large efficiency gains while preserving support for higher-precision quantized networks.